A railroad flatcar is travelling to the right at speed of 13 m/s relative to an observer on the ground. Someone is riding a motor scooter on the flatcar. What is the velocity of motor scooter relative to the flatcar if its relative velocity to the observer on the ground is (a) 18 m/s to the right? (b) 3 m/s to the left? (c) zero?


1
Expert's answer
2021-07-20T15:27:34-0400

Given:

"v=13\:\rm m/s"

(a)

"u=18-13=5\:\rm m/s \quad (right)"


(b)


"u=18+3=21\:\rm m/s \quad (left)"

(c)

"u=18+0=18 \:\rm m/s \quad (left)"